what is the amount of mole of sodiumtrioxocarbonate5 in 5.3g of the compound Na-23,C-12,O-16
1
Expert's answer
2012-10-31T08:17:08-0400
n(Na2CO3)= 5.3/ 106 = 0.05 mol;
n(Na) = 0.05 * 2 = 0.1 mol